# Augmented Physics: A Machine Learning-Powered Tool for Creating Interactive Physics Simulations from Static Diagrams

### v1

@@ 4.4.1. Kinematics Simulation
For our kinematics simulations, we utilize MatterJS222https://brm.io/matter-js/, a widely recognized JavaScript library for 2D Newtonian physics simulations. The images that users extract are transformed into 2D polygons that reflect their actual shapes. The polygons are subsequently integrated into the physics engine as rigid bodies, with their segmented images acting as sprites. These sprites receive user-defined properties, designating them as either static or dynamic bodies, where the former remains stationary and the latter is affected by simulation factors like gravity and time. Additionally, authors have the option to include a spring, a line, or identify an extracted object accordingly. The alignment of the segmented images and generated polygons is achieved by matching the bounding boxes of each object, ensuring that rigid bodies are accurately overlaid on the diagram.
@@ 4.4.2. Animated Diagrams
Our animation pipeline also uses the Segment-Anything model to extract the user-specified path from a diagram, offering a more effective solution than traditional line extraction methods. Utilizing the model’s ability to process both positive and negative prompts, users can easily identify their chosen path by clicking on it and marking it as a path. This feature allows for the addition of extra path points through further clicks on the image and the exclusion of undesired segments by identifying them as “negative” points, thereby improving the precision of path selection and the overall accuracy of path extraction. After isolating the path segment as an image mask, we apply skeletonization and thinning techniques with OpenCV and skimage to refine the mask to a 1-pixel width, effectively removing any extraneous artifacts or noise. The result is a clear set of points defining the intended path, ready to be animated. For the animation execution, we utilize the GSAP.js animation library333https://gsap.com/, animating the object along the determined path and integrating additional animation parameters, such as speed and direction, as required.
@@ 4.4.4. Circuit Simulation
We developed a custom circuit simulator designed to operate within a web browser, incorporating principles of circuit theory, such as Kirchhoff’s laws. Utilizing a support vector machine classifier, our system identifies and segments resistor, capacitor, and battery symbols within circuit diagrams. Further refinement is achieved through contour detection, which isolates lines and discriminates them based on their orientation. By identifying junctions within the diagram, the system automatically links the bounding boxes of detected resistors or voltage sources to lines, symbolizing wires. The circuit is represented using a simple array structure, which is transmitted to the frontend through the Firebase real-time database and subsequently visualized through simulation in a web browser setting.
@@ 4.5.1. Method
We evaluated the accuracy and versatility of our pipeline through technical evaluations. We first gathered six different physics textbooks covering topics such as kinematics, optics, circuit theory, and magnetism. From each textbook, we randomly selected 10 pages containing diagrams relevant to each simulation category (kinematics, circuits, optics, and animation), resulting in a total of 200 diagrams for our sample dataset. We applied our detection pipeline across these diagrams for each simulator category. For object segmentation, we simply select objects with mouse interaction. For line segmentation (for animated diagrams), we employ four points, two positive and two negative prompts, to segment the line. After that, multiple authors manually review the results by looking at the generated outcomes due to the absence of a standardized and automated way to check the results, guided by a rubric described below. The complete list of pages and figures evaluated with our system will be provided in the supplementary materials. Our analysis focused on measuring the error rate in various components of the pipeline.
@@ 4.5.2. Results
Table 1 presents a summary of our technical evaluation results. The success rates for the different components of the simulation are as follows: kinematics at 64%, optics at 44%, circuits at 15%, and animation at 66%.
Kinematics, optics, and animations depend on semi-automatic segmentation. Notably, object segmentation demonstrated a high success rate with 86% for both. This high success rate for segmentation contributes significantly to the relative success of kinematics, optics, and animation components. Specifically, in kinematics, the success rates for polygon generation and placement are 72% and 70%, respectively, indicating effective conversion into physics-simulatable bodies with proper segmentation. However, challenges arise in kinematics simulations due to limitations in supporting certain features (6%), such as rotational motion, body specific gravity, unsupported objects like ropes, and issues with simulating curved surfaces smoothly. Additionally, we noted that 74% of the diagrams just required minor adjustments, like modifications to simulation parameters, to achieve accurate simulation results. The success rate without any authoring and modification process was at 40%. Animation and optics were also consistent with the number, but we observed that the line segmentation success rate was lower, despite using the same Segment Anything technique.
On the other hand, circuit simulations utilize a custom automated recognition model, leading to a notably lower success rate of 15% for circuits. This lower rate is primarily attributed to the lack of support for various symbols, with 53% of diagrams containing unsupported elements such as capacitors, inductors, LEDs, and motors. Furthermore, our SVM-based automatic circuit symbol recognition system demonstrated a mere 10% success rate, hindered by an insufficient training dataset. Our method for circuit connection and line recognition achieved a 45% success rate. Overall, circuit simulations could perform well given the symbols are correctly recognized.
@@
| Supported Symbols | 47% | |
| Symbol Recognition | 15% | |
| Simulation | 15% |
@@ 5.1.1. Method
We recruited twelve participants, all of whom possessed at least high school-level knowledge in physics. Each session lasted approximately 40 minutes, and participants were compensated with 15 CAD for their time. After obtaining consent and introducing them to the study, we guided the participants through a demo walkthrough of the system, which included a circuit example (Figure 9). Once the participants had familiarized themselves with the system, they were allowed to interact with four prepared examples and features, including augmented experiments, animated diagrams, parameter visualization, and bi-directional manipulatives. All participants used the same textbook pages and physics diagrams, ensuring that everyone experienced the same set of features with a standardized set of examples. In our sessions, we utilized two optics examples, one on circuits, one on pendulums, one on kinematics with slope, and one animation featuring a solar system, as described in Section 4. We employed a talk-aloud methodology for the study, encouraging participants to verbalize their feedback as they interacted with the system.
@@ 5.2.2. Results
Experts generally conveyed that our system could help them create personalised simulations for their students and gave us invaluable feedback about our tool and future implementation. Below we summarize the expert feedback:
@@
Possible Features for Future Implementation
All experts wanted more customization in our system, advocating for more control over the scenario setup and the ability to add or duplicate custom objects for richer demonstrations. While seeking these enhancements, they emphasize the need to maintain simplicity to avoid a steep learning curve. They appreciated our system’s simplicity and said they they would want to be able to change parameters only when they want to. E6 and E11 also suggests incorporating visual aids such as trajectory paths or frame-by-frame illustrations to address observational challenges, like estimating speeds or remembering sequences. These possible features may further enhance the system’s educational value by making it more informative and accessible, linking theoretical concepts with visual evidence to deepen comprehension of physical principles.
@@ 6. Limitations and Future Work
Handling False Positives and Giving Teachers More Controls
To enable non-technical users to create physics augmentations, we used computer vision techniques such as Segment Anything and OpenCV to simplify the authoring process. Augmented Physics allows users to create dynamic and interactive simulations overlaid on physics textbooks on a touch based device. Although Augmented Physics supports authoring simulations for a variety of physics concepts, there are several domains Augmented Physics has not been developed to handle (that is, the corresponding simulator has not been built). More complex or abstract illustrations maybe too cluttered for our system to work well. Occasionally, Augmented Physics also fails to correctly detect objects from the image from its listed use cases (like resistors or lines). While participants valued the diversity of use cases we did build, some expressed a wish to use Augmented Physics on a wider range of physics concepts and demanded more sophisticated control over them. A commercially viable version of Augmented Physics would entail full usability in all topics of physics instruction. So, future implementations should include simulators for a variety of concepts in physics like molecular dynamics etc. Likewise, users should be afforded a larger set of controls to interact with simulated content. A fully deployed system can benefit from a plugin based system where teachers can port in their own simulators with which other can create dynamic and interactive embedded simulations to foster a more personalised teaching.
@@
In the future, we want to explore the use of AR interfaces beyond web-based interfaces. Furthermore, expanding beyond the mobile interface and introducing the system into an augmented reality (AR) headset environment could enhance user engagement. Although space constraints prevented direct diagram augmentation on the page, advancements in AR hardware may render this approach feasible. We are interested in developing augmented learning experiences on mixed reality devices, such as headsets. While current limitations in the precision of Augmented Physics preclude its immediate deployment on AR headsets, future refinements will open avenues for for immersive learning experiences beyond mobile and computer-based interfaces.
Deploying the System on a Large Scale
Our user study included expert interviews with experienced physics students, and interviews with non-experts. Although these participants highlighted the potential for using Augmented Physics in physics education, deploying Augmented Physics in classroom settings would better gauge it’s capabilities. In the future, we should examine how instructors can incorporate the features of Augmented Physics with their own teaching materials to enhance their teaching methods. We should also evaluate the deployment with novice physics learners to explore how they interact with the system to aid their learning on a large scale.
@@ 7. Conclusion
We introduced Augmented Physics, a machine learning-integrated tool for transforming static physics diagrams into interactive simulations. Using Segment-Anything and OpenCV, our system semi-automatically extracts content from diagrams, enabling non-technical users to create personalized, interactive, and animated explanations without programming. Findings from two user studies suggest that Augmented Physics supports more engaging and personalized learning experiences. Future work includes using our interaction techniques to other the domains covered in physics, investigating the potential for classroom use through in-the-wild deployments and lastly, exploring the mixed-reality modality to enhance physics education.

### v2

@@ 4.4.1. Kinematics Simulation
For our kinematics simulations, we utilize MatterJS222https://brm.io/matter-js/, a popular JavaScript library for 2D Newtonian physics simulations. The images that users extract are transformed into 2D polygons that reflect their actual shapes. The polygons are subsequently integrated into the physics engine as rigid bodies, with their segmented images acting as sprites. These objects receive user-defined properties, designating them as either static or dynamic bodies, where the former remains stationary and the latter is affected by simulation factors like gravity and time. Additionally, authors have the option to include a spring, a line, or identify an extracted object accordingly. The alignment of the segmented images and generated polygons is achieved by matching the bounding boxes of each object, ensuring that rigid bodies are accurately overlaid on the diagram.
@@ 4.4.2. Animated Diagrams
Our animation pipeline also uses the Segment-Anything model to extract the user-specified path from a diagram, offering a more effective solution than traditional line extraction methods. Utilizing the model’s ability to process both positive and negative prompts, users can easily identify their chosen path by clicking on it and marking it as a path. This feature allows for the addition of extra path points through further clicks on the image and the exclusion of undesired segments by identifying them as “negative” points, thereby improving the precision of path selection and the overall accuracy of path extraction. After isolating the path segment as an image mask, we apply skeletonization and thinning techniques with OpenCV and skimage to refine the mask to a 1-pixel width, effectively removing any extraneous artifacts and noise. The result is a clear set of points defining the intended path. The user can select multiple paths this way and assign them to multiple objects (extracted by SAM) in the diagram. For the animation execution, we utilize the GSAP.js animation library333https://gsap.com/, animating the object along the determined path and integrating additional animation parameters, such as speed and direction, editable by the user.
@@ 4.4.4. Circuit Simulation
We developed a custom circuit simulator designed to operate within a web browser, incorporating principles of circuit theory, such as Kirchhoff’s laws. Utilizing the Gemini Multimodal Vision Model (gemini-1.5-pro), our system identifies and segments resistor, capacitor, and battery symbols within circuit diagrams by detecting and extracting bounding boxes. Contour detection is then applied on the image, which isolates lines and discriminates them based on their orientation. By identifying junctions within the diagram, the system automatically links the bounding boxes of detected resistors or voltage sources to lines, symbolizing wires. The circuit is represented using a simple array structure, which is transmitted to the web interface through Firebase real-time database and subsequently visualized.
@@ 4.5.1. Method
We evaluated the accuracy and versatility of our pipeline through technical evaluations. We first gathered six different physics textbooks covering topics such as kinematics, optics, circuit theory, and magnetism. From each textbook, we randomly selected 10 pages containing diagrams relevant to each simulation category (kinematics, circuits, optics, and animation), resulting in a total of 200 diagrams for our sample dataset. We applied our detection pipeline across these diagrams for each simulator category. For object segmentation, we simply select objects via mouse interaction. For line segmentation (for animated diagrams), we employ four points, two positive and two negative prompts, to segment the line. After that, multiple authors manually review the results by looking at the generated outcomes due to the absence of a standardized and automated way to check the results, guided by a rubric described below. The complete list of pages and figures evaluated with our system will be provided in the supplementary materials. Our analysis focused on measuring the error rate in various components of the pipeline.
@@ 4.5.2. Results
Table 1 presents a summary of our technical evaluation results. The success rates for the different components of the simulation are as follows: kinematics at 64%, optics at 44%, circuits at 40% (62% with minor edits), and animation at 66%.
Kinematics, Optics, and Animation work through semi-automatic segmentation. Notably, object segmentation demonstrated a high success rate with 86%. This high success rate for segmentation contributes significantly to the relative success of kinematics, optics, and animation components. Specifically, in kinematics, the success rates for polygon generation and placement are 72% and 70%, respectively, indicating effective conversion into physics-simulatable bodies with proper segmentation. However, challenges arise in kinematics simulations due to limitations in supporting certain features (6%), such as rotational motion, body specific gravity, unsupported objects like ropes, and issues with simulating curved surfaces smoothly. Additionally, we noted that 74% of the diagrams just required minor adjustments, like modifications to simulation parameters, to achieve accurate simulation results. The success rate without any authoring and modification process was at 40%. Animation and optics were also consistent with the number, but we observed that the line segmentation success rate was lower, despite using the same Segment Anything technique. For optics in particular, simulation failures comes from diagrams our simulator does not support, such as those with multiple lenses (detecting two as one), prisms, new lens types (like an eye), etc.
Our Circuit simulation pipeline utilizes a line detection method to localize and identify wires in conjunction with the Gemini model to detect symbols. The line detection success rate is on the lower side at 45% leading to an overall success rate of simulation at 62% with minor connection edits. Without any edits the pipeline is successful 40% of the time. The main reason for this is because of overlapping or crossing wires or artifacts in the diagrams. On the other hand, symbol recognition is reasonably accurate at 72%. The symbols missed by the Gemini model were mainly due to the error in the bounding boxes returned, which did not align with the symbol.
@@
| Symbol Recognition | 72% | |
| Simulation | 40% | |
| Simulation with Minor Edits | 62% |
@@ 5.1.1. Method
We recruited 12 participants, all of whom possessed at least high school-level knowledge in physics. Each session lasted approximately 40 minutes, and participants were compensated with 15 CAD for their time. After obtaining consent and introducing them to the study, we guided the participants through a demo walkthrough of the system, which included a circuit example (Figure 9). Once the participants had familiarized themselves with the system, they were allowed to interact with four prepared examples and features, including augmented experiments, animated diagrams, parameter visualization, and bi-directional binding. All participants used the same textbook pages and physics diagrams, ensuring that everyone experienced the same set of features with a standardized set of examples. In our sessions, we utilized two optics examples, one on circuits, one on pendulums, one on kinematics with slope, and one animation featuring a solar system, as described in Section 4. We employed a talk-aloud methodology for the study, encouraging participants to verbalize their feedback as they interacted with the system.
@@ 5.1.2. Results
A horizontal stacked bar chart of the responses that respondents gave to our user study questionnaire. The length of each colored rectangle corresponds to the number of participants choosing a given response. Dark red denotes “strongly disagree”, orange denotes “disagree”, light orange denotes “somewhat disagree”, white denotes “neutral”, light blue denotes “somewhat agree”, blue denotes “agree” and dark blue denotes “strongly agree”. Response counts for each question are: Augmented Experiments help me in intuition building (strongly disagree=1, somewhat agree=1, agree=3, strongly agree=6) Animated Diagrams help me in intuition building (neutral=1, somewhat agree=2, agree=2, strongly agree=6) Parameter Visualization helps me in intuition building (agree=2, strongly agree=9) Bidirectional Manipulatives help me in intuition building (somewhat agree=1, agree=2, strongly agree=8)
@@ 5.2.2. Results
Experts generally conveyed that our system could help them create personalized simulations for their students and gave us invaluable feedback about our tool and future implementation. Below we summarize the expert feedback:
@@ 6. Limitations and Future Work
Deploying on a Large Scale
Despite the support for various physics concepts, the system struggles with complex or abstract illustrations and occasionally fails to detect certain objects. Participants appreciated the diversity of use cases but desired more sophisticated control and broader applicability. Future implementations can include simulators for a wider range of physics topics, like molecular dynamics, and offer more control over the simulations. Experts advocated for more customization, such as control over scenario setup and the ability to add or duplicate custom objects for richer demonstrations while maintaining simplicity to avoid a steep learning curve. E6 and E11 also suggested incorporating visual aids like trajectory paths or frame-by-frame illustrations to address observational challenges, like estimating speeds or remembering sequences. A plugin-based system could allow teachers to integrate custom simulators for broader applicability. Moreover, large-scale classroom deployment should be examined, evaluating how instructors and novice learners interact with the system to enhance learning on a larger scale.
@@
Expanding beyond the mobile interface and introducing the system into an augmented reality (AR) headset environment could enhance user engagement by providing a more immersive experience. While current limitations in the precision of Augmented Physics preclude its immediate deployment on AR headsets, future refinements will open avenues for immersive learning experiences beyond mobile and computer-based interfaces.
AI-Assisted Learning
Although we employ large language models (LLMs) in our pipeline for tasks such as detecting circuit symbols, recommending simulation types, and auto-filling parameter values, there is immense unexplored potential. Previous works have shown great promise in integrating intelligent recommendations into workflows such as online meetings (Xia et al., 2023b), active reading (Chen et al., 2022), and research (Fok et al., 2023). The next step could be intelligent physics tutoring. Integrating multimodal LLMs throughout our pipeline can enhance the authoring process. For instance, these models can automatically detect the context of diagrams, extract static and dynamic parts of images, and create simulations in real-time. Importantly, the system would still allow teachers to edit and reorganize simulations if the LLM makes errors or if the educator prefers to demonstrate a simplified version. Additionally, multimodal LLMs can function as teaching guides or tutoring agents, leading students through concepts step-by-step by creating simpler simulations and progressively increasing complexity. Exploring the efficacy of intelligent physics tutoring can be a promising future direction, integrating our extraction pipeline and augmentation strategies with LLMs in unique ways.
@@ 7. Conclusion
We introduced Augmented Physics, a machine learning-integrated tool for transforming static physics diagrams into interactive simulations. Using Segment-Anything, OpenCV, and MLLMs, our system semi-automatically extracts content from diagrams, enabling non-technical users to create personalized, interactive, and animated explanations without programming. Findings from two user studies suggest that our system supports more engaging and personalized learning experiences. Future work includes expanding to broader domains covered in physics, investigating the potential for classroom use through in-the-wild deployments and lastly, exploring the mixed-reality modality to enhance physics education.